HIGH RESOLUTION LASER SPECTROSCOPY OF SrOCH_{3}

\underline{D. FORTHOMME}, C. LINTON, D. W. TOKARYK, Centre for Laser, Atomic, and Molecular Sciences and Physics Department, 8 Bailey Dr., University of New Brunswick, P.O. Box 4400, Fredericton, NB, Canada E3B 5A3; A .G. ADAM, A. D. GRANGER, L. E. DOWNIE, W. S. HOPKINS, Centre for Laser, Atomic, and Molecular Sciences and Chemistry Department, 30 Dineen Dr., University of New Brunswick, P.O. Box 4400, Fredericton, NB, Canada E3B 5A3.

The Ã^{2}E−X̃^{2}A_{1} transition of SrOCH_{3} was first studied at high resolution by O’Brien et al.^{a} using a Broida oven. However lines with typically J≤20 were not observed and congestion prevented them from resolving transitions from the Ω = 1/2 component of the upper state. We have studied the B̃^{2}A_{1} − X̃^{2}A_{1} and the Ã^{2}E − X̃^{2}A_{1} transitions of SrOCH_{3} in a laser ablation molecular jet source, where jet cooling and low Doppler widths greatly simplified the spectra. An optical-optical double resonance technique facilitated definite assignments in a number of the transitions observed. Our analysis of the Ã^{2}E −X̃^{2}A_{1} transition was straightforward, but a perturbation was observed in the B̃^{2}A_{1} K′ = 1 F_{2} levels. A satisfactory fit was achieved for the B̃^{2}A_{1} − X̃^{2}A_{1} transition when a separate B parameter was used to fit the perturbed levels.

aL. C. O’Brien, C. R. Brazier and P. F. Bernath, J. Mol. Spectrosc. 130 (1988) 33-45